a) Find the eigenvalues and the eigenvectors of the 2x2 matrix:

[4   2]
[3 -1]

b) Solve the initial value problem:

dx/dt = 4x + 2y
dy/dt = 3x - y

with x(0) = 0, y(0) = 7
Answer #1

a) LA 3 lo-del=0 - 1 4 -2 1-0 = o (4-2) (-1-d) -G Hi=5 -2 d=5 [A-51] x=0 2 3 -6 = 2y V= 13] d- [A+QI]x=0 [s G 2 (31-6 o o 3 C[513]3-18 C = -> -4/ let ý = 3=) 01= -1 VE 3 5 -2 eigenvalves are are re [13]. [6 eigeove c-lorrs. CScanned with CamScannerb) 5- ge X(L) २- RA + Gear 3 X(E) = un २ce 5- + 39 २ ce X(0) [१] %3D CSScanned with CamScannerca 9t32 a 7 RC + Sss = IA -le = 2= 2 - 1 양야 Sect) = Qe ae 엇 LED = e XCE= 한 있e -de Re RL 5tt geal 약 7 beat e CScanned with Camwe first find eigenvalues and eigenvectors of A then we find solution x(t).then we apply initial condition to find constants.